OpenStudy (anonymous):

A team one 80% of its games. If it played 85 games, how many games did it win?

OpenStudy (anonymous):

85games * 80% wins = 85*.8 = 68 wins

OpenStudy (anonymous):

how do you calculate that exactly? im having trouble with percentages haha(: thanks for replying!!

OpenStudy (anonymous):

80% = 80/100 = 4/5 = .8 .8 * 85 = 68 :D

OpenStudy (anonymous):

10% of 85 games is 8.5. To get 80% multiply 8.5 by 8 to get 68